OpenStudy (socerlife66):

Which two values of x are roots of the polynomial below? x2 - 11x + 15

Vocaloid (vocaloid):

is there a list of answer choices?

Vocaloid (vocaloid):

use the quadratic formula

OpenStudy (campbell_st):

well looking at the problem you'll need the general quadratic formula to find them. the alternative is to take you choices, substitute the, and see which ones give a zero value.. hope it helps

jhonyy9 (jhonyy9):

x^2 -11x +15 = 0 D = 121 -4*15 = 121 -60 = 61 x_1,2 = (11 +/- sqrt61)/2 hope this will help you